### v1

# Agent²: An Agent-Generates-Agent Framework for Reinforcement Learning Automation

Abstract
Reinforcement learning agent development traditionally requires extensive expertise and lengthy iterations, often resulting in high failure rates and limited accessibility. This paper introduces Agent²—a novel agent-generates-agent framework that achieves fully automated RL agent design through intelligent LLM-driven generation. The system autonomously transforms natural language task descriptions and environment code into comprehensive, high-performance reinforcement learning solutions without human intervention.
Agent² features a revolutionary dual-agent architecture. The Generator Agent serves as an autonomous AI designer that analyzes tasks and generates executable RL agents, while the Target Agent is the resulting automatically generated RL agent. The framework decomposes RL development into two distinct stages: MDP modeling and algorithmic optimization, enabling more targeted and effective agent generation. Built on the Model Context Protocol, Agent² provides a unified framework that standardizes intelligent agent creation across diverse environments and algorithms, while incorporating adaptive training management and intelligent feedback analysis for continuous improvement.
Extensive experiments on a wide range of benchmarks—including MuJoCo, MetaDrive, MPE, and SMAC—demonstrate that Agent² consistently outperforms manually designed solutions across all tasks, achieving up to 55% performance improvement and substantial gains on average. By enabling truly end-to-end, closed-loop automation, this work establishes a new paradigm in which intelligent agents design and optimize other agents, marking a fundamental breakthrough for automated AI systems.
Introduction
Reinforcement learning (RL) has achieved remarkable success in diverse domains such as robotics, games, and autonomous systems. However, the process of developing high-performing RL agents remains notoriously complex, requiring substantial domain expertise, intricate environment engineering, careful algorithm selection, and tedious trial-and-error tuning (Li 2017). This manual, expertise-driven workflow has become a critical bottleneck to the practical adoption and democratization of RL, resulting in high entry barriers, limited scalability, and suboptimal research efficiency.
Recent advances in large language models (LLMs) have demonstrated unprecedented capabilities in autonomous reasoning and code synthesis (Achiam et al. 2023). These advancements make it possible to automate not just component-level tasks, but the entire RL agent development pipeline without human intervention.
In this work, we present Agent2, an agent-generates-agent framework that achieves end-to-end automation of RL agent design and optimization using LLMs. Agent2 introduces a novel dual-agent architecture: the Generator Agent acts as an autonomous AI designer capable of analyzing and generating all components required for RL agent; the resulting Target Agent is then constructed from these components and interacts with the environment for training and evaluation.
A key innovation of Agent2 is its intelligent two-stage decomposition of the RL development process. First, Agent2 convert high-level task descriptions into formal MDP representations, capturing environment dynamics, rewards, and constraints. Next, Agent2 automatically selects algorithms, designs neural architectures, and tunes hyperparameters, guided by LLM reasoning and empirical feedback. Integrated adaptive training and feedback analysis further enable robust error correction and iterative performance improvement through closed-loop interaction with the environment.
The remainder of this paper is organized as follows. We review related work, introduce our methodology, describe experimental details and results, and conclude with final remarks.
Related Work
Over the past few years, a wide range of general RL frameworks have been developed, such as RLlib (Liang et al. 2018), Stable Baselines3 (Raffin et al. 2021), Tianshou (Weng et al. 2022), and Xuance (Liu et al. 2023). These frameworks provide standardized implementations of algorithms, unified experiment management, and convenient interfaces for environment integration and distributed training. While these systems have greatly facilitated RL research and application, developing RL agents still heavily depends on expert knowledge and manual intervention. Currently, AutoML (He, Zhao, and Chu 2021) has seen rapid development and become relatively mature, but automated reinforcement learning (AutoRL) remains more complex and is still in its early stages. The ARLO framework (Mussi et al. 2023) systematically modularizes the AutoRL pipeline and supports automated selection and optimization of most stage. However, its automation is limited, as key stages such as feature engineering and data preparation often depend on hand-crafted methods and expert knowledge. Recently, LLMs have enabled new opportunities for RL automation by intelligently generating or optimizing key components of the RL pipeline (Cao et al. 2024; Sun, Huang, and Pompili 2024).
For MDP modeling, some studies have explored automating the design of state representations, reward functions, and action spaces. For state representation, YOLO-MARL (Zhuang et al. 2024) and LERO (Wei et al. 2025) use LLMs to infer high-level or hidden context in multi-agent environment, helping to address limitations from partial observability. In single-agent scenarios, LESR (Wang et al. 2024) and ExplorLLM (Ma et al. 2024a) use LLMs to autonomously generate or reformulate task-related state representations, improving learning efficiency and accelerating training. Reward function design has also been extensively studied in both single-agent and multi-agent RL. In particular, single-agent studies have addressed the challenge of sparse rewards by developing automated reward shaping methods, especially for tasks that require complex robotic control (Ma et al. 2024b; Song et al. 2023) or operate in high-level environments such as Minecraft (Li et al. 2024). In multi-agent settings, research mainly focuses on solving the credit assignment problem for effective reward distribution (Nagpal et al. 2025; Lin et al. 2025; He et al. 2025). For action spaces, recent methods use LLM-generated action masking or suboptimal policies to dynamically constrain and guide RL agents, incorporating user preferences and improving sample efficiency and policy adaptability (Zhao et al. 2025; Karine and Marlin 2025).
Beyond these components, the design of neural network architectures and the optimization of hyperparameters also play a crucial role in RL performance. Recent work has thus explored automating both aspects to further reduce manual effort. LLMatic (Nasir et al. 2024) combines LLM code generation with quality diversity optimization to discover diverse and effective architectures, while EvoPrompting (Chen, Dohan, and So 2023) uses LMs as adaptive operators in evolutionary NAS. In addition to network architecture design, recent studies have also leveraged LLMs to automate hyperparameter optimization in RL. Some studies use LLMs to suggest and iteratively refine hyperparameter configurations based on dataset and model descriptions (Liu, Gao, and Li 2024; Zhang et al. 2023). Additionally, SLLMBO (Mahammadli and Ertekin 2024) introduces a sequential LLM-based framework that adapts the search space dynamically and enhances optimization robustness. These methods highlight the potential of LLMs to design the network architecture and optimize the hyperparameter of RL algorithms.
However, most existing AutoRL approaches automate only a single stage of the RL pipeline, rather than enabling fully end-to-end, closed-loop agent generation. As a result, substantial expert intervention is still required to bridge the gap between task specification and practical agent deployment.
Methods
This section provides an overview of the Agent2 methodology, highlighting the automated workflow for RL agent generation. The overall process is illustrated in Figure 1. This section focuses on the Generator Agent, which serves as the core driver of automation in the Agent2 framework. The Generator Agent is responsible for analyzing tasks, structuring RL knowledge, and autonomously generating all components required for agent construction.
Figure 1: Agent2 processes natural language task descriptions, environment code, and constraints as input. Through problem analysis, it decomposes the task and assesses its complexity before invoking Model Context Protocol Services, which leverage LLMs for knowledge integration, natural language understanding, performance analysis, and code synthesis. The workflow then proceeds to MDP modeling—including state space, action space, and reward function design—followed by algorithmic optimization, where the agent autonomously selects algorithms, designs network architectures, and tunes hyperparameters. The generated components are assembled into the Target Agent, which is ready for training and evaluation. The entire process supports iterative refinement to enhance solution quality.
Task-to-MDP Mapping
Automatic Problem Analysis
A key prerequisite for automated RL agent generation is the precise understanding and formalization of the target problem. To address this, Agent2 begins the RL pipeline with a dedicated problem analysis stage, systematically bridging the gap between high-level user intent and formal RL formulation.
The Generator Agent leverages LLMs to process heterogeneous user input, typically including a natural language task description and, when available, environment code or specifications. The LLM integrates these inputs with its internal and, if configured, external knowledge bases, producing a structured representation of the RL problem. This analysis enables the system to: Identify learning objectives and reward structure from the task description; Infer state, action, and transition dynamics from the environment code; Extract explicit and implicit constraints, such as safety or operational limits; Anticipate challenges such as partial observability or delayed rewards.
[…]
$L_{\text{analysis}}=F_{\text{LLM}}(P_{\text{analysis}}(T_{\text{task}},T_{\text{env}},T_{c}))$ (1)
Here, $L_{\text{analysis}}$ is a structured analysis including formalized objectives, constraints, environment characteristics, and domain-specific challenges. This serves as the foundation for downstream MDP modeling and algorithm optimization, ensuring that all subsequent processes are both principled and context-aware.
MDP Modeling
The MDP modeling module automates the refinement of initial, suboptimal RL environments into effective MDPs for reinforcement learning. In practice, available simulators or environments may only provide a coarse or incomplete MDP specification, requiring expert effort to redesign the observation space, action space, or reward function. This module leverages LLMs to analyze and optimize these core components, minimizing manual intervention. While LLMs are capable of constructing MDPs from scratch, our approach focuses on the more practical and challenging scenario of automatically optimizing an existing but imperfect environment.
Formally, an MDP is defined as $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$, where $\mathcal{S}$ is the state (observation) space, $\mathcal{A}$ the action space, $\mathcal{P}$ the transition probability, $\mathcal{R}$ the reward function, and $\gamma$ the discount factor. Automated MDP modeling focuses on optimizing $\mathcal{S}$, $\mathcal{A}$, and $\mathcal{R}$, as these are central to agent-environment interaction.
The LLM examines environment code and task objectives to identify and refine relevant state variables, possibly applying feature selection, dimensionality reduction, or combining observations to improve learning efficiency. The refined observation space is expressed as $s^{\prime}=f_{\text{obs}}(s)\in\mathcal{S}^{\prime}$, where $f_{\text{obs}}$ denotes the learned or synthesized transformation function.
Similarly, the action space is optimized to balance expressiveness and simplicity, with the LLM recommending discretization, merging, splitting, or reparameterization as needed: $a^{\prime}=f_{\text{act}}(a)\in\mathcal{A}^{\prime}$, with $f_{\text{act}}$ constructed through automatic analysis and synthesis.
For reward design, the LLM constructs functions that accurately reflect task objectives and promote effective learning, addressing issues such as sparsity or bias. The reward at time $t$ is $r_{t}=f_{\text{rew}}(s_{t},a_{t},s_{t+1},\text{info})$, where $f_{\text{rew}}$ is the synthesized reward function.
Each component is encapsulated as a wrapper function, enabling seamless integration and algorithmic verification within the RL pipeline.
Adaptive Verification and Refinement
Automatically synthesized MDP components may have executability issues or lead to suboptimal learning due to ambiguities in code generation. To address this, we introduce an adaptive verification and refinement framework that integrates generated components into the RL pipeline and iteratively improves them via automated validation and feedback.
This module aims to ensure that all synthesized components are executable, and optimized for agent learning. Each component—observation, action, and reward functions ($f=\{f_{\text{obs}},f_{\text{act}},f_{\text{rew}}\}$)—is first integrated through standardized interfaces and checked by a verification operator $V$. If verification fails, an error feedback $e$ is generated and, together with the original RL analysis $L_{\text{analysis}}$ and component $f$, forms an adaptive prompt $P_{\text{error}}$. The LLM inference function $F_{\text{LLM}}$ then synthesizes a revised component $f^{*}$:
$\text{If}\;V(f)=\text{False},\;\text{then}\;f^{*}\leftarrow F_{\text{LLM}}\big{(}P_{\text{error}}(e,L_{\text{analysis}},f)\big{)}$ (2)
This process repeats until all components pass verification. Beyond static verification, the system continuously monitors the agent’s training dynamics and summarizes key performance indicators from TensorBoard data into a concise report, which serves as the performance feedback $\epsilon$ for further optimization. This feedback, together with $L_{\text{analysis}}$, $f$, and training history $\mathcal{H}$, to construct a prompt $P_{\text{perf}}$. The LLM then produces improved components:
$f^{*}\leftarrow F_{\text{LLM}}\big{(}P_{\text{perf}}(\epsilon,L_{\text{analysis}},f,\mathcal{H})\big{)}$ (3)
This closed-loop process ensures that MDP components are robust, reliable, and continually optimized, advancing the automation and practicality of LLM-based RL agent development.
The procedure is summarized in Algorithm 1.
Algorithm 1 Task to MDP Mapping Pipeline
0: Task description $T_{\text{task}}$, environment information $T_{\text{env}}$, additional constraints $T_{c}$, LLM inference function $F_{\text{LLM}}$
0: Verified and optimized MDP components $f=\{f_{\text{obs}},f_{\text{act}},f_{\text{rew}}\}$
1:// Automatic Problem Analysis
[…]
3:// MDP Modeling
[…]
5:// Adaptive Verification and Refinement
6: Initialize best components $f^{*}\leftarrow f^{(0)}$, best score $S^{*}_{\text{hist}}\leftarrow S^{(0)}$, training history $\mathcal{H}\leftarrow\{\}$
7:for $t=1$ to $T$ do
8:  Integrate $f^{(t)}$ into RL pipeline
9:  Apply verification operator $V$ to $f^{(t)}$
10:  if verification fails (error eethen
11:   $f^{(t+1)}\leftarrow F_{\text{LLM}}(P_{\text{error}}(e,L_{\text{analysis}},f^{(t)}))$
12:   continue
13:  end if
14:  Train agent with $f^{(t)}$, collect TensorBoard metrics, update $\mathcal{H}$
15:  Summarize key metrics into feedback $\epsilon^{(t)}$
16:  Obtain performance score $S^{(t)}$
17:  if $S^{(t)}>S^{*}_{\text{hist}}$ then
18:   $S^{*}_{\text{hist}}\leftarrow S^{(t)}$, $f^{*}\leftarrow f^{(t)}$
19:  end if
20:  if not converged then
21:   $f^{(t+1)}\leftarrow F_{\text{LLM}}(P_{\text{refine}}(\epsilon^{(t)},L_{\text{analysis}},f^{(t)},\mathcal{H}))$
22:  else
23:   break
24:  end if
25:end for
26:return  Final verified and optimized MDP components $f=\{f_{\text{obs}},f_{\text{act}},f_{\text{rew}}\}$
Algorithmic Optimization
Upon completing the task-specific MDP design, Agent² automatically performs algorithm selection, network architecture design, training hyperparameter optimization, and integrated configuration with iterative refinement. The details of these four stages are presented in the following subsections.
Algorithm Selection
Algorithm selection aims to match the constructed MDP with an RL algorithm suitable for the environment’s characteristics and task requirements. For example, value-based methods like DQN are appropriate for discrete action spaces, whereas policy gradient methods such as PPO and SAC are better for continuous control. Choosing the right algorithm is crucial for effective and stable policy learning.
Agent2 uses LLMs to adaptively select algorithms based on comprehensive information, including the structured RL analysis $L_{\text{analysis}}$, environment code $T_{\text{env}}$, and current MDP components $f=\{f_{\text{obs}},f_{\text{act}},f_{\text{rew}}\}$. Given a set of candidates $G$, the LLM identifies the most suitable algorithm $g^{*}\in G$ and provides a brief rationale $L_{\text{algo}}$. Formally, the prompt is $P_{\text{alg}}(L_{\text{analysis}},f,T_{\text{env}})$, and the LLM inference function performs:
$g^{*},\ L_{\text{algo}}=F_{\text{LLM}}\big{(}P_{\text{alg}}(G,f,L_{\text{analysis}},T_{\text{env}})\big{)}$ (4)
Network Architecture Design
Effective reinforcement learning requires neural network architectures tailored to both the environment and the selected algorithm $g^{*}$. Agent2 integrates information from observation and action space descriptions, the algorithm, structured reference architectures (e.g., $J_{\text{net}}$ in JSON), and statistical summaries from training history ($\mathcal{H}$). Using these structured prompts, the LLM critiques and refines previous designs, leveraging both domain expertise and empirical results to suggest architectural improvements.
Formally, the network design process is defined as:
$N^{*},\ L_{\text{net}}=F_{\text{LLM}}\Big{(}P_{\text{net}}(g^{*},f_{\text{obs}},f_{\text{act}},J_{\text{net}},\mathcal{H})\Big{)}$ (5)
where $N^{*}$ is the recommended architecture and $L_{\text{net}}$ explains the design. This automated approach enables Agent2 to generate adaptive network architectures for reliable RL training across diverse environments.
Hyperparameter Optimization
Training performance in RL depends heavily on careful hyperparameter selection and adjustment, such as learning rate, discount factor, batch size, and regularization. Agent2 uses an LLM-driven process to systematically initialize and refine hyperparameters for the current environment, algorithm $g^{*}$, and network architecture $N^{*}$.
The Generator Agent consolidates structured information from the MDP model ($f_{\text{obs}}$, $f_{\text{act}}$, $f_{\text{rew}}$), network $N^{*}$, selected algorithm $g^{*}$, reference configurations $J_{\text{hp}}$, and training history $\mathcal{H}$. Prompt engineering guides the LLM to assess current settings, identify bottlenecks, and propose incremental, interpretable adjustments prioritized for training stability and performance.
Formally, hyperparameter optimization is defined as:
$Hp^{*},\ L_{\text{hp}}=F_{\text{LLM}}\Big{(}P_{\text{hp}}(g^{*},N^{*},f_{\text{obs}},f_{\text{act}},f_{\text{rew}},J_{\text{hp}},\mathcal{H})\Big{)}$ (6)
where $Hp^{*}$ is the optimized hyperparameter set and $L_{\text{hp}}$ provides concise justifications for each change.
Configuration Integration and Refinement
After algorithm selection, network architecture design, and hyperparameter optimization, Agent2 automatically integrates the selected algorithm $g^{*}$, network $N^{*}$, and optimized hyperparameters $Hp^{*}$ into a unified configuration $\mathcal{C}$, which is exported in standardized YAML format for compatibility and reproducibility.
Agent2 also supports adaptive refinement: after deployment, it continuously monitors training performance through TensorBoard data collection and processing. When needed, the LLM is leveraged to incrementally update $\mathcal{C}$ based on observed results and historical data $\mathcal{H}$, forming a closed-loop process that continuously enhances agent performance and robustness.
The complete algorithmic optimization pipeline is summarized in algorithm 2.
Algorithm 2 Algorithmic Optimization Pipeline
0: MDP components $f=\{f_{\text{obs}},f_{\text{act}},f_{\text{rew}}\}$, problem analysis $L_{\text{analysis}}$, environment code $T_{\text{env}}$, LLM inference function $F_{\text{LLM}}$
0: Optimized RL configuration $\mathcal{C}^{*}$
1:// Algorithm Selection
[…]
3:// Network Architecture Design
[…]
5:// Hyperparameter Optimization
[…]
7:// Configuration Integration
8: Merge $g^{*}$, $N^{*}$, and $Hp^{*}$ into configuration $\mathcal{C}$
9:// Adaptive Refinement
10: Initialize best configuration $\mathcal{C}^{*}\leftarrow\mathcal{C}^{(0)}$, best score $S^{*}_{\text{hist}}\leftarrow S^{(0)}$, training history $\mathcal{H}\leftarrow\{\}$
11:for $t=1$ to $T$ do
12:  Integrate $\mathcal{C}^{(t)}$ into RL training pipeline
13:  Train agent with $\mathcal{C}^{(t)}$, collect TensorBoard metrics, update $\mathcal{H}$
14:  Summarize key metrics into feedback $\epsilon^{(t)}$
15:  Obtain performance score $S^{(t)}$
16:  if $S^{(t)}>S^{*}_{\text{hist}}$ then
17:   $S^{*}_{\text{hist}}\leftarrow S^{(t)}$, $\mathcal{C}^{*}\leftarrow\mathcal{C}^{(t)}$
18:  end if
19:  if not converged then
20:   $\mathcal{C}^{(t+1)}\leftarrow F_{\text{LLM}}(P_{\text{refine}}(\epsilon^{(t)},L_{\text{analysis}},\mathcal{C}^{(t)},\mathcal{H}))$
21:  else
22:   break
23:  end if
24:end for
25:return  Final best configuration $\mathcal{C}^{*}$
Experiments
We empirically validate Agent2 on diverse single-agent and multi-agent RL benchmarks, comparing its performance to original baselines and analyzing the impact of each component.
Experimental Setup
Environments.
Single-agent environments include classic MuJoCo continuous control tasks—Ant, Humanoid, Hopper, and Walker. Each task requires an agent to control a simulated robot for stable and efficient locomotion: Ant involves quadruped movement, Humanoid requires balancing a high-dimensional biped, Walker2d targets bipedal walking, and Hopper focuses on single-legged hopping. We also include the more challenging MetaDrive (Li et al. 2022), a large-scale autonomous driving simulator where agents must safely navigate diverse and dynamic traffic scenarios using rich sensory observations.
Multi-agent environments comprise the Multi-Agent Particle Environment (MPE) (Lowe et al. 2017) with two scenarios: Simple Spread, where agents cooperate to cover all landmarks with minimal collision, and Simple Reference, which combines movement and communication as agents must infer and reach assigned targets. Additionally, we use the StarCraft Multi-Agent Challenge (SMAC) (Samvelyan et al. 2019), evaluating on tasks of increasing difficulty: 8m (homogeneous Marines team), 2s3z (heterogeneous Stalker and Zealot units), and 1c3s5z, a large-scale, heterogeneous team requiring advanced coordination and tactical planning.
Baselines.
We select widely used reinforcement learning algorithms as baselines to ensure the generality and credibility of our comparisons. For single-agent experiments, we employ PPO, SAC, and TD3 on the MuJoCo environments, and PPO and SAC on MetaDrive. These algorithms are among the most popular and well-established methods in the continuous control domain, and are commonly used as benchmarks in major RL frameworks.
For all multi-agent scenarios, we adopt MAPPO, a classic algorithm that has demonstrated strong performance in many benchmarks such as MPE and SMAC. MAPPO is widely recognized for its stability and scalability in cooperative tasks and is supported by mainstream RL codebases.
As baselines, we use the original MDP model of all scenarios with the default hyperparameter settings of all algorithms, as implemented in the Xuance framework(Liu et al. 2023). To ensure fair and consistent comparison, we report the cumulative episode reward as the primary evaluation metric for all environments. For StarCraft scenarios, we additionally report the task success rate to better reflect the performance on complex cooperative tasks.
Experimental Details.
All experiments are conducted on a workstation equipped with a 22-core AMD EPYC 7402 CPU, a NVIDIA RTX 4090 GPU, and 108 GB of RAM. Throughout the automated configuration and optimization process, Claude-Sonnet-3.7 is utilized as the underlying LLMs. Notably, Claude-Sonnet-3.7 has demonstrated strong capabilities in problem analysis and code generation, making it a robust choice for our experiments.
For each experiment, we set the total environment interaction steps as follows: 1mm steps for MuJoCo tasks, 400kk steps for MetaDrive, 1mm steps for MPE scenarios, and 400kk, 1mm, or 2mm steps for the different SMAC tasks. The automated optimization pipeline uses 5 rounds of MDP generation followed by 5 rounds of algorithm configuration for MuJoCo, MetaDrive, and MPE environments, while 3 rounds of each are used for SMAC tasks.
Experimental Results
MuJoCo.
Across a suite of MuJoCo continuous control tasks, Agent2 consistently achieves higher cumulative rewards and better learning dynamics compared to the baseline. In Ant-v4, PPO improves from 3280.51 to 3831.86, TD3 from 3853.84 to 5981.39, and SAC from 3499.76 to 4372.05 (see Fig. 2(a)). The training curves (see Fig. 3(a), 3(b), 3(c)) further show substantial improvements in sample efficiency and final performance, especially for TD3 and SAC. After the baseline quickly converges, the episode rewards of Agent2 continue to increase significantly in the later stages of training.
In the more challenging Humanoid-v4 environment, Agent2 demonstrates exceptional robustness. Notably, the baseline performance of TD3 is extremely poor (354.79) due to an unsuitable default configuration, while Agent2 dramatically lifts the score to 5425.46. PPO and SAC also show clear improvements, with final scores rising from 917.64 to 1085.85 and from 4682.83 to 6787.94, respectively (see Fig. 2(b)). The corresponding training curves (see Fig. 3(d), 3(e), 3(f)) illustrate: while Agent2 with PPO shows a modest improvement after the baseline stabilizes, Agent2 with SAC and TD3 display large further gains after the baseline converges, highlighting the strong late-stage optimization of Agent2.
For Hopper-v4 and Walker2d-v4, Agent2 also brings observable gains across all algorithms (see Fig. 2(c), Fig. 2(d)).
Figure 2: Performance comparison on MuJoCo benchmarks.
Figure 3: Training curve comparisons on MuJoCo benchmarks.
MetaDrive.
On the MetaDrive autonomous driving benchmark, Agent2 consistently outperforms the baseline for both PPO and SAC. With automated optimization, the cumulative reward for PPO increases from 245.34 to 277.54, and for SAC from 178.16 to 259.76. As shown in Fig. 4(a), Agent2 achieves clear overall performance gains. The training curves in Fig. 5(a), 5(b) show that, Agent2 with PPO achieves a moderate increase in the final convergence level, while Agent2 with SAC substantially accelerates early training progress and rapidly surpasses the baseline from the beginning. These results highlight the effectiveness of automated optimization even in complex, real-world-like single-agent environments.
MPE.
On MPE cooperative tasks, Agent2 yields clear improvements in team performance. For Simple Spread, the average episode reward improves from -19.73 to -16.31. For Simple Reference, the reward rises from -19.61 to -19.00. The overall performance comparison is shown in Fig. 4(b), and the training curves is shown in Fig.5(c), 5(d).
Figure 4: Performance comparison on MetaDrive and MPE benchmarks.
Figure 5: Training curve comparisons on MetaDrive and MPE benchmarks.
SMAC.
On challenging SMAC scenarios, Agent2 consistently improves the task win rate over the baseline. In the 8m scenario, win rate increases from 0.77 to 0.94; in 2s3z, from 0.82 to 0.85; and in the more difficult 1c3s5z, from 0.17 to 0.23. Figure 6 provides an overview of the performance gains across all scenarios. The training curves in Fig.7(a),7(b),7(c) show that in the 8m scenario, Agent2 achieves both faster convergence and substantially higher final episode rewards compared to the baseline. In the more challenging 1s3c5z scenario, Agent2 initially matches the baseline, but surpasses it in both cumulative rewards and final performance as training progresses. Moreover, the win rate curves in Fig.7(d),  7(e),  7(f) reflect similar trends, with Agent2 achieving higher win rates, especially in complex environments where the improvements are even more pronounced. These results highlight the effectiveness of automated optimization for multi-agent tactical cooperation in challenging scenarios.
Figure 6: Performance comparison on the StarCraft-II benchmark.
Figure 7: Training curve comparisons on StarCraft II benchmark.
Ablation Analysis
To rigorously demonstrate the effectiveness of our two-stage pipeline, we quantitatively compare the baseline (default manual configuration), after Task-to-MDP Mapping (Stage 1), and after subsequent Algorithmic Optimization (Stage 2). As summarized in Figure 8, most environments and algorithms show clear improvements: Task-to-MDP Mapping improves performance in 83% of scenarios, and Algorithmic Optimization delivers additional gains in 67% of cases.
After Stage 1, In Ant-v4 with TD3, the average episode reward jumps from 3853.8 (baseline) to 5981.4, an impressive 55% improvement. In the challenging SMAC-1c3s5z scenario, the win rate increases from 0.17 to 0.21 (a 25% relative gain), and in MetaDrive with SAC, the reward rises from 178.2 to 219.6 (a 23% increase). For most environments, Task-to-MDP Mapping alone enables the agent to outperform the baseline by a large margin, demonstrating the effectiveness of automatic MDP adaptation.
Second, Algorithmic Optimization further boosts performance, especially in more challenging tasks or with sensitive algorithms. For instance, in Ant-v4 with PPO, Stage 2 brings an additional 13% improvement over Stage 1 (3831.9 vs. 3385.4). In Humanoid-v4 with SAC, performance improves by another 11.6% after Stage 2 (6787.9 vs. 6081.2). Similarly, in MetaDrive with SAC, the reward increases by an additional 18% (259.8 vs. 219.6), and in SMAC-1c3s5z, the win rate continues to climb from 0.21 to 0.23 (a further 10% relative gain).
These findings highlight the complementary nature of our pipeline: Task-to-MDP Mapping accurately constructs a specialized training MDP from natural language task descriptions, enabling rapid adaptation of the agent to the environment structure, while Algorithmic Optimization fine-tunes learning for maximum effectiveness. Together, these stages enable robust and fully automated generation of high-performing RL agents across a diverse range of single-agent and multi-agent benchmarks.
Figure 8: Performance improvement comparison across Task-to-MDP Mapping and Algorithmic Optimization.
Conclusion
In this work, we present Agent2, a fully automated, LLM-driven agent-generates-agent framework that significantly advances the automation, accessibility, and performance of RL development. By introducing a novel dual-agent architecture and a two-stage pipeline that separates MDP modeling from algorithmic optimization, Agent2 is able to autonomously translate high-level task descriptions and raw environment code into high-performing RL agents—entirely without human intervention.
Extensive experiments across single-agent and multi-agent benchmarks demonstrate that Agent2 consistently outperforms manually designed baselines, achieving performance improvements across all evaluated environments and algorithms, which validates both the effectiveness and generality of the proposed approach. Our ablation analysis further highlights the complementary impact of automated environment adaptation and iterative algorithmic refinement. Notably, Agent2 achieves performance improvement in 83% of task-algorithm pairs after automated MDP adaptation, and a further gain in 67% of cases following algorithmic optimization.
Ultimately, Agent2 establishes a new paradigm in which intelligent agents can autonomously design, optimize, and deploy other agents, thereby breaking through traditional barriers of RL development and setting the stage for the next generation of automated AI systems.