### v2

# Agent²: An Agent-Generates-Agent Framework for Reinforcement Learning Automation

Abstract
Reinforcement learning (RL) agent development traditionally requires substantial expertise and iterative effort, often leading to high failure rates and limited accessibility. This paper introduces Agent2, an LLM-driven agent-generates-agent framework for fully automated RL agent design. Agent2 autonomously translates natural language task descriptions and environment code into executable RL solutions without human intervention.
The framework adopts a dual-agent architecture: a Generator Agent that analyzes tasks and designs agents, and a Target Agent that is automatically generated and executed. To better support automation, RL development is decomposed into two stages—MDP modeling and algorithmic optimization—facilitating targeted and effective agent generation. Built on the Model Context Protocol, Agent2 provides a unified framework for standardized agent creation across diverse environments and algorithms, incorporating adaptive training management and intelligent feedback analysis for continuous refinement.
Extensive experiments on benchmarks including MuJoCo, MetaDrive, MPE, and SMAC show that Agent2 outperforms manually designed baselines across all tasks, achieving up to 55% performance improvement with consistent average gains. By enabling a closed-loop, end-to-end automation pipeline, this work advances a new paradigm in which agents can design and optimize other agents, underscoring the potential of agent-generates-agent systems for automated AI development.
1 Introduction
Reinforcement learning (RL) has achieved remarkable success in diverse domains such as robotics, games, and autonomous systems. However, the process of developing high-performing RL agents remains notoriously complex, requiring substantial domain expertise, intricate environment engineering, careful algorithm selection, and tedious trial-and-error tuning (Li, 2017). This manual, expertise-driven workflow has become a major barrier to the practical adoption and broader accessibility of RL, leading to high entry costs, limited scalability, and reduced research efficiency. Recent advances in large language models (LLMs) have demonstrated unprecedented capabilities in autonomous reasoning and code synthesis (Achiam et al., 2023). These advancements make it possible to automate not just component-level tasks, but the entire RL agent development pipeline without human intervention.
In this work, we present Agent2, an agent-generates-agent framework that introduces a novel dual-agent architecture. The Generator Agent serves as an autonomous AI designer, capable of analyzing and producing all necessary components for an RL agent. Based on these components, the Target Agent is constructed and subsequently interacts with the environment for training and evaluation.
A key innovation of Agent2 lies in its end-to-end automated pipeline for RL agent generation. Instead of merely automating isolated components, Agent2 achieves unified and functionalized development of each module, efficient information exchange across modules, and integrated training frameworks. Moreover, Agent2 incorporates mechanisms for detecting and resolving training anomalies, enabling iterative evolution and continual refinement. This comprehensive design transforms RL development into a adaptive and progressively improving pipeline. The remainder of this paper is organized as follows. Section 2 reviews related work, Section 3 presents our methodology, Section 4 reports experimental results, and Section 5 concludes the paper.
2 Related Work
Over the past few years, a wide range of general RL frameworks have been developed, such as RLlib (Liang et al., 2018), Tianshou (Weng et al., 2022), and Xuance (Liu et al., 2023). These frameworks provide standardized implementations of algorithms, unified experiment management, and convenient interfaces for environment integration and distributed training. While these frameworks have greatly facilitated RL research and application, developing RL agents still heavily depends on expert knowledge and manual intervention. LLMs have enabled new opportunities for RL automation by intelligently generating or optimizing key components of the RL pipeline (Cao et al., 2024; Sun et al., 2024).
We contend that the RL automation can be broadly categorized into two complementary dimensions: MDP modeling and algorithmic optimization. For MDP modeling, some studies have explored automating the design of state representations, reward functions, and action spaces. LESR (Wang et al., 2024) and ExplorLLM (Ma et al., 2024a) use LLMs to autonomously generate or reformulate task-related state representations, improving learning efficiency and accelerating training. YOLO-MARL (Zhuang et al., 2024) and LERO (Wei et al., 2025) use LLMs to infer high-level or hidden context in multi-agent environment, helping to address limitations from partial observability. Some studies of reward function design have addressed the challenge of sparse rewards by developing automated reward shaping methods, especially for tasks that require complex robotic control (Ma et al., 2024b; Song et al., 2023) or operate in high-level environments such as Minecraft (Li et al., 2024). In multi-agent settings, research mainly focuses on solving the credit assignment problem for effective reward distribution (Nagpal et al., 2025; Lin et al., 2025; He et al., 2025). For action spaces, recent methods use LLM-generated action masking or suboptimal policies to dynamically constrain and guide RL agents, incorporating user preferences and improving sample efficiency and policy adaptability (Zhao et al., 2025; Karine & Marlin, 2025).
Research on its algorithmic optimization mainly follows the AutoML, which has seen rapid development and become relatively mature (He et al., 2021). Algorithmic optimization focusing on two directions: neural network architecture design and hyperparameter optimization. LLMatic (Nasir et al., 2024) combines LLM code generation with quality diversity optimization to discover diverse and effective architectures, while EvoPrompting (Chen et al., 2023) uses LMs as adaptive operators in evolutionary NAS. For hyperparameter optimization, some studies use LLMs to suggest and iteratively refine hyperparameter configurations based on dataset and model descriptions (Liu et al., 2024; Zhang et al., 2023; Mussi et al., 2023). Additionally, SLLMBO (Mahammadli & Ertekin, 2024) introduces a sequential LLM-based framework that adapts the search space dynamically and enhances optimization robustness. These methods highlight the potential of LLMs to design the network architecture and optimize the hyperparameter of RL algorithms.
However, most existing AutoRL approaches automate only a single stage of the RL pipeline, rather than enabling fully end-to-end agent generation, which requires the coordination of multiple modules as well as the handling of their interactions and debugging. This poses a far more challenging problem, which we address in the following section.
3 Methods
This section provides an overview of the Agent2 methodology, highlighting the automated workflow for RL agent generation. The overall process is illustrated in Figure 1. This section focuses on the Generator Agent, which serves as the core driver of automation in the Agent2 framework. The Generator Agent is responsible for analyzing tasks, structuring RL knowledge, and autonomously generating all components required for agent construction.
Figure 1: The framework of Agent2 consists of three main stages. Firstly, Agent2 analyzes the problem using natural language task descriptions and environment code as inputs. Secondly, the framework proceeds to MDP modeling, including the design of the state space, action space, and reward function. Thirdly, the framework is followed by algorithmic optimization, where the agent autonomously selects appropriate algorithms, designs network architectures, and tunes hyperparameters. The entire framework operates in compliance with the Model Context Protocol (MCP), ensuring standardized integration of services. Finally, the generated components are assembled into the Target Agent, which is ready for training and evaluation. The entire process supports iterative refinement to enhance solution quality.
3.1 Task-to-MDP Mapping
3.1.1 Automatic Problem Analysis
A key prerequisite for automated RL agent generation is the precise understanding and formalization of the target problem. The Generator Agent leverages LLMs to process heterogeneous user input with its knowledge bases, producing a structured representation of the RL problem. This analysis enables the system to: Identify learning objectives and reward structure from the task description; Infer state, action, and transition dynamics from the environment code; Extract explicit and implicit constraints, such as safety or operational limits; Anticipate challenges such as partial observability or sparse rewards.
[…]
$L_{\text{analysis}}=F_{\text{LLM}}(P_{\text{analysis}}(T_{\text{task}},T_{\text{env}},T_{c}))$ (1)
Here, $L_{\text{analysis}}$ is a structured analysis including formalized objectives, constraints, environment characteristics, and domain-specific challenges. This serves as the foundation for downstream MDP modeling and algorithm optimization, ensuring that all subsequent processes are both principled and context-aware.
3.1.2 MDP Modeling
The MDP modeling module automates the refinement of initial RL environments into effective MDPs for reinforcement learning. In practice, available simulators or environments may only provide a coarse or incomplete MDP specification, requiring expert effort to redesign the observation space, action space, or reward function. This module leverages LLMs to analyze and optimize these core components, minimizing manual intervention. While LLMs are capable of constructing MDPs from scratch, our approach focuses on the more practical and challenging scenario of automatically optimizing an existing but imperfect environment.
Formally, an MDP is defined as $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$, where $\mathcal{S}$ is the state (observation) space, $\mathcal{A}$ the action space, $\mathcal{P}$ the transition probability, $\mathcal{R}$ the reward function, and $\gamma$ the discount factor. Automated MDP modeling focuses on optimizing $\mathcal{S}$, $\mathcal{A}$, and $\mathcal{R}$, as these are central to agent-environment interaction.
The LLM examines environment code and Problem Analysis to identify and refine relevant state variables, possibly applying feature selection, dimensionality reduction, or combining observations to improve learning efficiency. The refined observation space is expressed as $s^{\prime}=f_{\text{obs}}(s)\in\mathcal{S}^{\prime}$, where $f_{\text{obs}}$ denotes the learned or synthesized transformation function.
Similarly, the action space is optimized to balance expressiveness and simplicity, with the LLM recommending merging, splitting, or reparameterization as needed: $a^{\prime}=f_{\text{act}}(a)\in\mathcal{A}^{\prime}$, with $f_{\text{act}}$ constructed through automatic analysis and synthesis.
For reward design, the LLM constructs functions that accurately reflect task objectives and promote effective learning, addressing issues such as sparsity or bias. The reward at time $t$ is $r_{t}=f_{\text{rew}}(s_{t},a_{t},s_{t+1},\text{info})$, where $f_{\text{rew}}$ is the synthesized reward function.
Each component is encapsulated as a wrapper function, enabling seamless integration and algorithmic verification within the RL framework.
3.1.3 Adaptive Verification and Refinement
Automatically synthesized MDP components may have executability issues or lead to suboptimal learning due to ambiguities in code generation. To address this, we introduce an adaptive verification and refinement framework that integrates generated components into the RL pipeline and iteratively improves them via automated validation and feedback.
Each component—observation, action, and reward functions ($f=\{f_{\text{obs}},f_{\text{act}},f_{\text{rew}}\}$)—is first integrated through standardized interfaces and checked by a verification operator $V$. If verification fails, an error feedback $e$ is generated and, together with the original RL analysis $L_{\text{analysis}}$ and component $f$, forms an adaptive prompt $P_{\text{error}}$. The LLM inference function $F_{\text{LLM}}$ then synthesizes a revised component $f^{*}$.
Beyond static verification, the system continuously monitors the agent’s training dynamics and summarizes key performance indicators from TensorBoard data into a concise report, which serves as the performance feedback $\epsilon$ for further optimization. This feedback, together with $L_{\text{analysis}}$, $f$, and training history $\mathcal{H}$, to construct a prompt $P_{\text{perf}}$. The LLM then produces improved components. This closed-loop process ensures that MDP components are robust, reliable, and continually optimized, advancing the automation and practicality of LLM-based RL agent development.
The procedure is summarized in Algorithm 1.
Algorithm 1 Adaptive Verification and Refinement of MDP Components
0: Initial MDP components $f^{(0)}=\{f_{\text{obs}},f_{\text{act}},f_{\text{rew}}\}$, LLM inference function $F_{\text{LLM}}$, RL analysis $L_{\text{analysis}}$, maximum iterations $T$
0: Verified and optimized MDP components $f^{*}$
1: Initialize best components $f^{*}\leftarrow f^{(0)}$, best score $S^{*}_{\text{hist}}\leftarrow S^{(0)}$, training history $\mathcal{H}\leftarrow\{\}$
2:for $t=1$ to $T$ do
3:  Integrate $f^{(t)}$ into RL pipeline
4:  Apply verification operator $V$ to $f^{(t)}$
5:  if $V(f^{(t)})=\text{False}$ with error $e$ then
6:   $f^{(t+1)}\leftarrow F_{\text{LLM}}(P_{\text{error}}(e,L_{\text{analysis}},f^{(t)}))$
7:   continue
8:  end if
9:  Train agent with $f^{(t)}$, collect TensorBoard metrics $\epsilon^{(t)}$, obtain score $S^{(t)}$, update $\mathcal{H}$
10:  if $S^{(t)}>S^{*}_{\text{hist}}$ then
11:   $S^{*}_{\text{hist}}\leftarrow S^{(t)}$, $f^{*}\leftarrow f^{(t)}$
12:  end if
13:  if not converged then
14:   $f^{(t+1)}\leftarrow F_{\text{LLM}}(P_{\text{perf}}(\epsilon^{(t)},L_{\text{analysis}},f^{(t)},\mathcal{H}))$
15:  else
16:   break
17:  end if
18:end for
19:return $f^{*}$
3.2 Algorithmic Optimization
Upon completing the task-specific MDP design, Agent² automatically performs algorithm selection, network architecture design, training hyperparameter optimization, and integrated configuration with iterative refinement. The details of these four stages are presented in the following subsections.
3.2.1 Algorithm Selection
Algorithm selection aims to match the constructed MDP with an RL algorithm suitable for the environment’s characteristics and task requirements. For example, value-based methods like DQN are appropriate for discrete action spaces, whereas policy gradient methods such as PPO and SAC are better for continuous control. Choosing the right algorithm is crucial for effective and stable policy learning.
Agent2 uses LLMs to adaptively select algorithms based on comprehensive information, including the structured RL analysis $L_{\text{analysis}}$, environment code $T_{\text{env}}$, current MDP components $f=\{f_{\text{obs}},f_{\text{act}},f_{\text{rew}}\}$, and training history $\mathcal{H}$. Given a set of candidates $G$, the LLM identifies the most suitable algorithm $g^{*}\in G$ and provides a brief rationale $L_{\text{algo}}$. Formally, the LLM inference function performs:
$g^{*},\ L_{\text{algo}}=F_{\text{LLM}}\big(P_{\text{alg}}(G,f,L_{\text{analysis}},T_{\text{env}},\mathcal{H})\big)$ (2)
3.2.2 Network Architecture Design
Effective reinforcement learning requires neural network architectures tailored to both the environment and the selected algorithm $g^{*}$. Agent2 integrates information from observation and action space descriptions, the algorithm, structured reference architectures (e.g., $J_{\text{net}}$ in JSON), and statistical summaries from training history ($\mathcal{H}$). Using these structured prompts, the LLM critiques and refines previous designs, leveraging both domain expertise and empirical results to suggest architectural improvements.
Formally, the network design process is defined as:
$N^{*},\ L_{\text{net}}=F_{\text{LLM}}\Big(P_{\text{net}}(g^{*},f_{\text{obs}},f_{\text{act}},J_{\text{net}},\mathcal{H})\Big)$ (3)
where $N^{*}$ is the recommended architecture and $L_{\text{net}}$ explains the design. This automated approach enables Agent2 to generate adaptive network architectures for reliable RL training across diverse environments.
3.2.3 Hyperparameter Optimization
Training performance in RL depends heavily on careful hyperparameter selection and adjustment, such as learning rate, discount factor, batch size, and regularization. Agent2 uses an LLM-driven process to systematically initialize and refine hyperparameters for the current environment, algorithm $g^{*}$, and network architecture $N^{*}$.
The Generator Agent consolidates structured information from the MDP model $f^{*}$, network $N^{*}$, selected algorithm $g^{*}$, reference configurations $J_{\text{hp}}$, and training history $\mathcal{H}$. Prompt engineering guides the LLM to assess current settings, identify bottlenecks, and propose incremental, interpretable adjustments prioritized for training stability and performance.
Formally, hyperparameter optimization is defined as:
$Hp^{*},\ L_{\text{hp}}=F_{\text{LLM}}\Big(P_{\text{hp}}(g^{*},N^{*},f^{*},J_{\text{hp}},\mathcal{H})\Big)$ (4)
where $Hp^{*}$ is the optimized hyperparameter set and $L_{\text{hp}}$ provides concise justifications for each change.
3.2.4 Configuration Integration and Refinement
After algorithm selection, network architecture design, and hyperparameter optimization, Agent2 automatically integrates the selected algorithm $g^{*}$, network $N^{*}$, and optimized hyperparameters $Hp^{*}$ into a unified configuration $\mathcal{C}$, which is exported in standardized YAML format for compatibility and reproducibility. For adaptive refinement, Agent2 monitors training performance via TensorBoard metrics and leverages the LLM to incrementally update $\mathcal{C}$ using observed results and historical data $\mathcal{H}$, forming a closed-loop process that enhances performance and robustness.
The pipeline of configuration integration and refinement is summarized in algorithm 2.
Algorithm 2 Algorithmic Configuration Integration and Refinement
0: Selected algorithm $g^{*}$, network architecture $N^{*}$, optimized hyperparameters $Hp^{*}$, optimized MDP components $f^{*}$, problem analysis $_{\text{analysis}}, LL$M inference function $F_{\text{LLM}}$
0: Final best configuration $\mathcal{C}^{*}$
1: Initialize best configuration $\mathcal{C}^{*}\leftarrow\mathcal{C}^{(0)}$, best score $S^{*}_{\text{hist}}\leftarrow S^{(0)}$, training history $\mathcal{H}\leftarrow\{\}$
2:for $t=1$ to $T$ do
3:  Merge $g^{*}$, $N^{*}$, and $Hp^{*}$ into configuration $\mathcal{C}$
4:  Train agent with $\mathcal{C}^{(t)}$, collect TensorBoard metrics $\epsilon^{(t)}$, obtain score $S^{(t)}$, update $\mathcal{H}$
5:  if $S^{(t)}>S^{*}_{\text{hist}}$ then
6:   $S^{*}_{\text{hist}}\leftarrow S^{(t)}$, $\mathcal{C}^{*}\leftarrow\mathcal{C}^{(t)}$
7:  end if
8:  if not converged then
9:   Execute equations (2), (3), (4) to adaptive refinement $g^{*}$, $N^{*}$, $Hp^{*}$
10:  else
11:   break
12:  end if
13:end for
14:return Final best configuration $\mathcal{C}^{*}$
4 Experiments
To comprehensively evaluate the effectiveness of Agent2, we conduct experiments in progressive stages. We begin with a comparison under the MuJoCo benchmark, where our framework is evaluated against several well-established RL libraries, for which benchmark results are publicly available in this setting. We then extend the evaluation to a broader set of single-agent and multi-agent environments to test the generality and robustness of our approach. Finally, we perform ablation studies to quantify the respective impact of the two stages: Task-to-MDP mapping and algorithmic optimization.
4.1 Experiment Setup
Environments. Single-agent environments include classic MuJoCo continuous control tasks—Ant, Humanoid, Hopper, and Walker2d. Each task requires an agent to control a simulated robot for stable and efficient locomotion. We also include the more challenging MetaDrive (Li et al., 2022), a large-scale autonomous driving simulator where agents must safely navigate diverse and dynamic traffic scenarios using rich sensory observations. Multi-agent environments include MPE (Lowe et al., 2017) with Simple Spread for cooperative landmark coverage and Simple Reference for communication-based target reaching, and SMAC (Samvelyan et al., 2019), which provides cooperative StarCraft II micromanagement tasks of varying scales and difficulties that require advanced coordination and tactical planning.
Baselines. We select widely used reinforcement learning algorithms as baselines to ensure the generality and credibility of our comparisons. For single-agent experiments, we employ PPO, SAC, and TD3 on the MuJoCo environments, and PPO and SAC on MetaDrive. For all multi-agent scenarios, we adopt MAPPO, a classic algorithm that has demonstrated strong performance in many benchmarks such as MPE and SMAC. As baselines, we use the original MDP model of all scenarios with the default hyperparameter settings of all algorithms, as implemented in the Xuance framework(Liu et al., 2023). In our framework, we employ Claude-Sonnet-3.7 as the LLM, which demonstrates stable performance in code generation. For more details about the experiment setup, please refer to Appendix A.
4.2 Experimental Results
MuJoCo benchmark comparison. Firstly, we benchmark Agent2 on a suite of MuJoCo tasks, directly comparing its performance with several widely used RL libraries. As shown in Table 1, since our framework is built on Xuance, we first compare directly against Xuance results. In almost all scenarios, Agent2 outperforms Xuance, with substantial improvements in the Ant (e.g., PPO: 3831.9 vs.2810.7; TD3: 5981.4 vs.4822.9) and Humanoid (e.g., SAC: 6788.0 vs.4682.8) tasks; the only exception occurs in Hopper with PPO, where the performance is comparable (3444.6 vs.3450.1). Notably, the performance of Humanoid with TD3 in XuanCe benchmark is extremely poor (547.8) due to an unsuitable default configuration, while Agent2 dramatically lifts the score to 5425.5. Compared to Tianshou, Agent2 demonstrates clear advantages in most cases, with only two instances—Ant with SAC and Walker2d with SAC—where its performance is marginally weaker. Against SpinningUp, Agent2 consistently outperforms across all reported benchmarks.
Table 1: The comparison between Agent2 with classic RL benchmarks in Mujoco environment
Scenario Algorithm XuanCe Tianshou SpinningUp Agent2
Ant PPO 2810.7 3258.4 650 3831.9
Ant TD3 4822.9 5116.4 3800 5981.4
Ant SAC 3499.7 5850.2 3980 4372.1
Humanoid PPO 705.5 787.1 - 1085.9
Humanoid TD3 547.8 5189.5 - 5425.5
Humanoid SAC 4682.8 5488.5 - 6788.0
Hopper PPO 3450.1 2609.3 1850 3444.6
Hopper TD3 3492.4 3472.2 3564.1 3570.1
Hopper SAC 3517.4 3542.2 3150 3576.5
Walker2d PPO 4318.6 3588.5 1230 4649.5
Walker2d TD3 4307.9 3982.4 4000 4844.4
Walker2d SAC 4730.5 5007.0 4250 4805.3
Generalization to broader environments. To assess generality and robustness beyond MuJoCo, we further evaluate Agent2 across a wider range of single-agent and multi-agent environments. Since these extended environments lack publicly available benchmark results, we compare Agent2 against the Xuance framework, ensuring fairness by adopting the same training budget and reporting the best performance achieved.
In Ant scenario of MuJoCo environment, the training curves (see Fig. 2(a), 2(b), 2(c)) show substantial improvements in sample efficiency and final performance, especially for TD3 and SAC. After the baseline quickly converges, the episode rewards of Agent2 continue to increase significantly in the later stages of training. In the more challenging Humanoid environment, The corresponding training curves (see Fig. 2(d), 2(e), 2(f)) illustrate: while Agent2 with PPO shows a modest improvement after the baseline stabilizes, Agent2 with SAC and TD3 display large further gains after the baseline converges, highlighting the strong late-stage optimization of Agent2.
(a) PPO-Ant
(b) SAC-Ant
(c) TD3-Ant
(d) PPO-Humanoid
(e) SAC-Humanoid
(f) TD3-Humanoid
Figure 2: Training curve comparisons on MuJoCo environments.
On the MetaDrive autonomous driving environment, Agent2 consistently outperforms the baseline for both PPO and SAC. The training curves in Fig. 3(a), 3(b) show that, Agent2 with PPO achieves a moderate increase in the final convergence level, while Agent2 with SAC substantially accelerates early training progress and rapidly surpasses the baseline from the beginning. These results highlight the effectiveness of automated optimization even in complex, real-world-like single-agent environments.
On MPE cooperative tasks, Agent2 brings modest yet consistent gains in team performance. In Simple Spread, the average episode reward improves from -19.73 to -16.31, while in Simple Reference, it increases from -19.61 to -19.00. The corresponding training curves (Fig. 3(c), 3(d)) show these steady improvements over the baseline.
(a) PPO-MetaDrive
(b) SAC-MetaDrive
(c) MAPPO-Spread
(d) MAPPO-Reference
Figure 3: Training curve comparisons on MetaDrive and MPE environments.
On challenging SMAC scenarios, Agent2 consistently improves over the baseline. The training curves in Fig.4(a), 4(b), 4(c) show that in the 8m scenario, Agent2 achieves both faster convergence and substantially higher final episode rewards compared to the baseline. In the more difficult 1c3s5z scenario, Agent2 initially matches the baseline, but gradually surpasses it in cumulative rewards and final performance as training progresses. The win rate curves (Fig.4(d), 4(e), 4(f)) reflect similar trends, with win rates rising from 0.77 to 0.94 in 8m, from 0.82 to 0.85 in 2s3z, and from 0.17 to 0.23 in 1c3s5z. Overall, the results indicate that Agent2 delivers consistent performance improvements, particularly in complex scenarios, thereby validating the effectiveness of automated optimization for multi-agent tactical cooperation.
(a) MAPPO-SC2-8m
(b) MAPPO-SC2-2s3z
(c) MAPPO-SC2-1c3s5z
(d) MAPPO-SC2-8m
(e) MAPPO-SC2-2s3z
(f) MAPPO-SC2-1c3s5z
Figure 4: Training curve comparisons on StarCraft-II environments.
Ablation Analysis. To rigorously demonstrate the effectiveness of our two-stage pipeline, we quantitatively compare the baseline (default manual configuration), after Task-to-MDP Mapping (Stage 1), and after subsequent Algorithmic Optimization (Stage 2). As summarized in Figure 5, most environments and algorithms show clear improvements: Task-to-MDP Mapping improves performance in 83% of scenarios, and Algorithmic Optimization delivers additional gains in 67% of cases. We exclude the humanoid-TD3 experiment, as it exhibited abnormal behavior in the benchmark comparison.
Compared with the baseline, Stage 1 optimizes the task formulation by reconstructing the MDP, delivering substantial gains across domains. For example, in Ant-TD3 the reward increases by 55% (3853.8 → 5981.4), in MetaDrive-SAC by 23% (178.2 → 219.6), and in the challenging SMAC-1c3s5z scenario, the win rate improves from 0.17 to 0.21 (a 25% relative gain). Building upon this, Stage 2 further enhances performance by refining algorithmic configurations: Ant-PPO achieves an additional 13% improvement (3385.4 → 3831.9), Humanoid-SAC rises by 11.6% (6081.2 → 6787.9), MetaDrive-SAC gains another 18% (219.6 → 259.8), and SMAC-1c3s5z advances from 0.21 to 0.23 (a further 10% relative gain). Together, these stages enable robust and fully automated generation of high-performing RL agents across a diverse range of single-agent and multi-agent benchmarks. Detailed statistics can be found in Appendix B.
Figure 5: Performance improvement across Task-to-MDP Mapping and Algorithmic Optimization.
5 Conclusion
In this work, we present Agent2, a fully automated, LLM-driven agent-generates-agent framework that advances the automation, accessibility, and performance of RL development. By introducing a novel dual-agent architecture and two-stage pipeline separating MDP modeling from algorithmic optimization, Agent2 autonomously translates high-level task descriptions and raw environment code into high-performing RL agents without human intervention.
Extensive experiments on single-agent and multi-agent benchmarks show that Agent2 outperforms the majority of manually designed baselines across diverse environments and algorithms, demonstrating both the effectiveness and generality of the approach. Ablation analysis further highlights the complementary roles of automated environment adaptation and iterative algorithmic refinement: Agent2 boosts performance in 83% of task-algorithm pairs after MDP adaptation, with further gains in 67% of cases following optimization. Together, these findings highlight the potential of Agent2 for real-world deployment and point toward future extensions to increasingly complex tasks and domains.
Appendix A Experimental Details.
A.1 Environments.
For the MuJoCo suite, we consider four classic continuous control tasks. In Ant, the agent controls a four-legged ant robot and aims to move forward as fast and smoothly as possible, keeping the body stable and preventing falls. The Humanoid scenario is more challenging, requiring the agent to coordinate a high-dimensional humanoid robot to walk quickly and stably while maintaining balance. In Walker2d, the agent must control a two-legged bipedal robot to walk forward, applying torques to six joints to ensure stability. The Hopper task involves a single-legged robot that must hop forward rapidly and avoid falling, requiring fine control over three actuated joints.
MetaDrive serves as a large-scale autonomous driving simulator, where the agent must safely drive a vehicle through randomly generated roads and traffic scenarios. The objective is to reach the destination efficiently while avoiding collisions with other vehicles and obstacles. The observation space incorporates ego-vehicle states, navigation cues, and lidar-like sensor data, providing a comprehensive view of the driving environment.
The Multi-Agent Particle Environment (MPE) is used to evaluate multi-agent cooperation and communication. In the Simple Spread scenario, multiple agents must collaborate to cover all target landmarks on the map while minimizing collisions with each other. Each agent observes its own velocity and position, as well as information about landmarks and other agents. In the Simple Reference scenario, two agents are each tasked with approaching specific colored landmarks, but the assignment is communicated by the other agent. Agents must leverage both movement and communication actions to succeed, receiving both individual and global rewards.
For more complex cooperative tasks, we use the StarCraft Multi-Agent Challenge (SMAC) benchmark. In 8m, eight allied Marines must coordinate to defeat eight enemy Marines on a symmetric map with no obstacles, emphasizing position and focus fire. The 2s3z scenario is heterogeneous, featuring teams of two Stalkers and three Zealots each, which requires the agent to adapt strategies for mixed unit types. 1c3s5z is a large-scale heterogeneous battle with one Colossus, three Stalkers, and five Zealots per side, demanding advanced coordination and tactical planning due to unit diversity and increased team size.
A.2 Experimental Parameters
In our experiments, we conducted a thorough evaluation of the proposed framework across multiple environments and algorithms. All experiments are conducted on a workstation equipped with a 22-core AMD EPYC 7402 CPU, a NVIDIA RTX 4090 GPU, and 108 GB of RAM. Throughout the automated process, we employ Claude-Sonnet-3.7 as the underlying LLM. This model has shown stable performance in both problem analysis and code generation, and thus serves as a dependable backbone for our framework. Table 2 provides a summary of the model parameters used in our experiments.
Table 2: Model inference parameters.
Parameter Value Description
Context length 128K Maximum input tokens supported by the model
Max_tokens 1024 Maximum tokens generated in one response
Temperature 0.6 Controls randomness; lower = more deterministic
Top_p 0.7 Nucleus sampling cutoff; balances diversity
Top_k 50 Limits sampling to the top-$k$ probable tokens
To ensure reproducibility and provide a clear understanding of our experimental setup, we established standardized parameters for all evaluations. In terms of training settings, we carefully balanced training costs with performance outcomes to configure appropriate parameters that would demonstrate the effectiveness of our approach across diverse scenarios without excessive computational requirements. Table 3 presents the experimental configurations of each environment, including the training steps, the evaluation episodes, and the number of automated optimization iterations.
Table 3: Experimental settings for different environments
Environment Training Steps Evaluation episodes iterations per stage
MuJoCo 1M 50 5
MetaDrive 400K 50 5
MPE 1M 50 5
SMAC-8m 400K 50 3
SMAC-2s3z 1M 50 3
SMAC-1c3s5z 2M 50 3
A.3 LLM Prompt Design
To clearly describe the design workflow, we provide the complete LLM prompt used in our study as follows.
<System>
You are a professional reinforcement learning problem analysis expert.
Please analyze the given problem and output the analysis results in the following format:
# Task Objectives
[Describe specific optimization goals]
# Constraints
[List main limitations]
# Environment Characteristics
- Deterministic/Stochastic: [Explanation]
- Fully/Partially Observable: [Explanation]
- Single-agent/Multi-agent: [Explanation]
# Key Challenges
[List 2-3 main challenges]
<User>
[…]
Listing 1: Problem Analyzing Prompt
<System>
You are an expert RL observation space designer specializing in optimizing state representations.
Your task is to design the most effective observation space based on problem description and analysis, and implement an ObsWrapper function to transform original observation vectors into new observation vectors.
Design Objectives:
1. Improve agent learning efficiency through better state representation
2. Enhance feature extraction and information utilization
3. Normalize and scale values appropriately
4. Balance information richness with dimensionality
Pay attention to division by zero issues in floating-point scalar operations, avoid generating nan in calculations.
Please output strictly in the following format:
# Observation Variables
[List main observation variables and their meanings]
# Observation Space Definition of Single-agent
‘‘‘json
{"dim": dimension, "type": "continuous/discrete", "low": minimum_value, "high": maximum_value}
# Observation Space Definition of Multi-agent
‘‘‘json
"agent_0": {"dim": dimension, "type": "continuous/discrete", "low": minimum_value, "high": maximum_value},
"agent_1": {"dim": dimension, "type": "continuous/discrete", "low": minimum_value, "high": maximum_value},
# ObsWrapper Implementation of Single-agent
‘‘‘python
def custom_state_transform(state: np.ndarray) -> np.ndarray:
# Implement state transformation logic
return transformed_state
# ObsWrapper Implementation of Multi-agent
‘‘‘python
def custom_state_transform(state: Dict[str, np.ndarray]) -> Dict[str, np.ndarray]:
# Transform each agent separately
transformed_state = {}
for agent_id, agent_obs in state.items():
# Transformation logic
transformed_state[agent_id] = ...
return transformed_state
# Design Description
[Brief explanation of design ideas and considerations, including:
1. Key transformations applied and their purpose
2. How this design addresses the specific challenges of the environment
3. Expected impact on agent learning]
<User>
Task Environment: {env_name} - {scenario_name}
Problem Description: {problem_description}
Problem Analysis: {analysis_str}
Environment Code: {env_code}
default Observation Space: {default_observation_space}
Design Requirements:
1. Focus on extracting meaningful features from raw observations
2. Apply appropriate normalization to stabilize learning
3. Consider feature engineering that highlights task-relevant information
4. Maintain numerical stability in all transformations
5. If the observation space is discrete, the space dim equals the number of discrete observations
Listing 2: Observation Space Designing Prompt
<System>
You are an expert RL action space designer specializing in optimizing action representations.
Your task is to design the most effective action space based on problem description, analysis and observation space,
and implement an ActionWrapper function that maps the designed action vector to the environments original action space.
Design Objectives:
1. Improve agent learning efficiency through better action representation
2. Create intuitive action mappings that facilitate policy learning
3. Balance expressiveness with simplicity
4. Ensure smooth transitions between action spaces
Pay attention to division by zero issues in floating-point scalar operations, avoid generating nan in calculations.
Please output strictly in the following format:
# Action Variables
[List main action variables and their meanings]
# Action Space Definition of Single_agent
‘‘‘json
{"dim": dimension, "type": "continuous/discrete", "low": minimum_value, "high": maximum_value}
# Action Space Definition of Multi-agent
‘‘‘json
"agent_0": {"dim": dimension, "type": "continuous/discrete", "low": minimum_value, "high": maximum_value},
"agent_1": {"dim": dimension, "type": "continuous/discrete", "low": minimum_value, "high": maximum_value},
# ActionWrapper Implementation of Single_agent
‘‘‘python
def custom_action_transform(custom_action: np.ndarray) -> np.ndarray:
# Implement action transformation logic
return gym_action
# ActionWrapper Implementation of Multi-agent
‘‘‘python
def custom_action_transform(custom_action: Dict[str, np.ndarray]) -> Dict[str, np.ndarray]:
# Transform each agent separately
gym_action = {}
for agent_id, agent_action in custom_action.items():
# Transformation logic
gym_action[agent_id] = ...
return gym_action
# Design Description
[Brief explanation of design ideas and considerations, including:
1. Key transformations applied and their purpose
2. How this design addresses the specific challenges of the environment
3. Expected impact on agent learning]
<User>
Task Environment: {env_name} - {scenario_name}
[…]
Design Requirements:
1. Compatibility with transformed observation space features
2. Smoothness and numerical stability of action transformation
3. Avoid action space being too complex leading to training difficulties
4. Create intuitive mappings between designed actions and environment actions
5. If the action space is discrete, the space dim equals the number of discrete actions
Listing 3: Action Space Designing Prompt
<System>
You are an expert RL reward function designer specializing in optimizing reward signals.
Your task is to design the most effective reward function based on problem description, analysis, environment code, observation space, and action space,
and implement a RewardWrapper function that calculates rewards based on custom current state, action, next state and environment information.
Design Objectives:
1. Create a reward signal that effectively guides learning toward desired behaviors
2. Balance immediate feedback with long-term goals
3. Avoid reward hacking and degenerate policies
4. Ensure numerical stability and proper scaling
Pay attention to division by zero issues in floating-point scalar operations, avoid generating nan in calculations.
Please output in the following format:
# Reward Calculation Logic
[Explain the main components and weights of reward calculation]
# RewardWrapper Implementation of Single_agent
‘‘‘python
def custom_reward_function(
custom_current_state: np.ndarray,
custom_action: np.ndarray,
custom_next_state: np.ndarray,
info: dict
) -> float:
# Implement reward calculation logic
return custom_reward
# RewardWrapper Implementation of Multi-agent
‘‘‘python
def custom_reward_function(
custom_current_state: Dict[str, np.ndarray],
custom_action: Dict[str, np.ndarray],
custom_next_state: Dict[str, np.ndarray],
info: dict
) -> Dict[str, float]:
# Calculate rewards for each agent separately
custom_reward = {}
for agent_id in custom_current_state:
# Reward calculation logic
custom_reward[agent_id] = ...
return custom_reward
# Design Description
[Brief explanation of design ideas and considerations, including:
1. Key reward components and their purpose
2. How this design addresses the specific challenges of the environment
3. Expected impact on agent learning]
<User>
Task Environment: {env_name} - {scenario_name}
Problem Description: {problem_description}
Problem Analysis: {analysis_str}
Environment Code: {env_code}
Designed Observation Space Code: {obs_code}
Implemented Observation Space: {obs_space}
Designed Action Space Code: {action_code}
Implemented Action Space: {action_space}
Design Requirements:
1. Input states have been transformed by ObsWrapper, please refer to ObsWrapper code to understand state structure
2. Prioritize using info dictionary to get environment information, avoid hard-coding extraction from state vectors
3. Ensure numerical stability of reward calculation, avoid nan or inf values
4. Create a reward structure that guides learning toward optimal behavior
5. Balance immediate feedback with long-term goals
Listing 4: Reward function Designing Prompt
<System>
You are an expert in reinforcement learning neural network architecture design.
Given the observation space and action space, please design the most suitable network architecture for the problem.
For each item, strictly choose only from the options listed in parentheses.
Your response should follow this format:
# Network Architecture and Parameters, provide detailed parameter configuration, including:
1. Layer Types and Dimensions (choose only from: Basic_MLP, Basic_Identical, Basic_CNN, Basic_RNN; specify dimensions)
2. Activation Functions (choose only from: relu, leaky_relu, tanh, sigmoid, softmax, elu)
3. Regularization Methods (choose only from: LayerNorm, BatchNorm, BatchNorm2d)
4. Other Special Configurations (if any; otherwise leave blank)
# Design Description
[Brief explanation of design ideas and considerations]
<User>
Task Environment: {env_name} - {scenario_name}
Algorithm: {algorithm}
Target Parameters for Optimization: {list(network_config.keys())}
Observation Space: {obs_design_str}
Action Space: {action_design_str}
Listing 5: Network Designing Prompt
<System>
You are an expert RL hyperparameter tuner specializing in optimizing specific training parameters.
Your job is to set appropriate training hyperparameters based on environment, algorithm, neural network design and reference configuration.
Tuning Objectives:
1. Improve the overall performance.
2. Enhance training stability.
3. Respect computational constraints.
4. Favor incremental improvements where possible.
Recommended Hyperparameter Tuning Priorities:
Learning Rate: A crucial factor for training stability and convergence.
Discount Factor (gamma): Balances the importance of immediate and future rewards. Common values range from 0.97 to 0.995.
GAE Lambda (gae_lambda): Adjusts the bias-variance trade-off in advantage estimation. Typical values are between 0.92 and 0.97.
Clip Range: Helps maintain stable policy updates. Consider values within [0.1, 0.3].
Batch Size and Update Frequency: Parameters such as horizon size, number of epochs, and minibatch count should be tuned to balance learning speed and stability.
Value Function and Entropy Coefficients: Tuning the value loss coefficient (e.g., 0.1-0.5) and entropy coefficient (e.g., 0.0-0.01) can improve value estimation and exploration, respectively.
Gradient Clipping: Prevents large, destabilizing updates. Experiment with different gradient clipping norms, such as 0.5 or 1.0.
Observation and Reward Normalization: Enabling and tuning normalization can enhance training stability and generalization.
Output Format:
‘‘‘yaml
parameter_name: value
# Reason: [problem identified] - [why change helps] - [expected outcome]
Requirements:
- Start hyperparameter searches with the learning rate and clip range, then proceed to gamma, gae_lambda, and batch/update parameters. Fine-tune normalization and coefficient settings in later stages for best results.
- Suggest changes to only relevant parameters (max 5).
- Use incremental changes (less than 30%) unless fixing critical stability issues.
- Ensure values remain within safe, specified ranges.
- Consider parameter interactions and algorithm-specific constraints.
- Give concise, clear reasoning for each suggestion.
<User>
Task Environment: {env_name} - {scenario_name}
Algorithm: {algorithm}
Network Architecture: {network_design}
Target Parameters for Optimization: {list(algorithm_config.keys())}
Observation Space: {obs_design_str}
Action Space: {action_design_str}
Please analyze the current configuration and determine if any of the target parameters need optimization.
Listing 6: Hyperparameters Designing Prompt
Appendix B Experimental Results
To demonstrate the effectiveness of our two-stage pipeline, we quantitatively compare the baseline (default manual configuration), after Task-to-MDP Mapping (Stage 1), and after subsequent Algorithmic Optimization (Stage 2). The detailed results are summarized in Table 4.
Table 4: Performance comparison on 2 stages
Environment Scenario Algorithm Baseline Stage1 Stage2
MuJoCo Ant PPO 3280.51 3385.38 3831.86
MuJoCo Ant TD3 3853.84 5981.39 5981.39
MuJoCo Ant SAC 3499.76 3960.73 4372.05
MuJoCo Humanoid PPO 917.64 1085.85 1085.85
MuJoCo Humanoid TD3 354.79 5425.46 5425.46
MuJoCo Humanoid SAC 4682.83 6081.16 6787.94
MuJoCo Hopper PPO 3304.03 3314.88 3444.56
MuJoCo Hopper TD3 3486.29 3570.13 3570.13
MuJoCo Hopper SAC 3517.41 3517.41 3576.54
MuJoCo Walker2d PPO 4236.68 4649.51 4649.51
MuJoCo Walker2d TD3 4513.16 4844.38 4844.38
MuJoCo Walker2d SAC 4730.53 4730.53 4805.28
MetaDrive metadrive PPO 245.34 264.57 277.54
MetaDrive metadrive SAC 178.16 219.60 259.76
MPE simple_spread_v3 MAPPO -19.73 -16.63 -16.31
MPE simple_reference_v3 MAPPO -19.61 -19.00 -19.00
SC2 8m MAPPO 0.77 0.92 0.94
SC2 2s3z MAPPO 0.82 0.84 0.85
SC2 1c3s5z MAPPO 0.17 0.21 0.23